7:12 Instagram Done 4 of 13 PARTPRORF REPORT THE NUMERICAL VALUES AND SYM ANSWER EACH ONESTION WITHIN THE SPACE PROVIDED NSWEplease answer and show work
Answer #1

1). Loss on each share = share price at expiration - exercise price = 45-40 = 5 per share

So, total loss = (number of shares*-5) + option premium = -500 + 4 = -496

2). FV = 1,000; PV = -1,025; PMT = 6%*FV = 6%*1,000 = 60; N = 10, solve for RATE. YTM = 5.67%

3). Fund X beta = sum of weighted betas

Formula I/Total w*b
Stock Investment (I) Weight (w) Beta (b)
A 200000 10.0% 1.5 0.15
B 300000 15.0% -0.5 -0.075
C 500000 25.0% 1.25 0.3125
D 1000000 50.0% 0.75 0.375
Total 2000000 100.0% 0.7625

Fund X beta = 0.7625

Using CAPM, required return = risk-free rate + beta*(market return - risk-free rate)

= 7% + 0.7625*(13.25%-7%) = 11.77%